# Onsager vortex clusters on a sphere

### v1

@@ IV.3.1 Sandwich states at high energies
$\frac{1}{4}\frac{1}{\rho}\frac{{\rm d}}{{\rm d}\rho}\left(\rho\frac{{\rm d}}{{\rm d}\rho}\psi\right)=-e^{-\beta\psi(\rho)+\gamma_{+}},$ (70)
@@
$\psi(\rho)$ $=-\frac{2}{\beta}\log\left(\frac{2}{\beta\rho^{2}e^{\gamma_{+}}-2}\right)$ (71)
with boundary condition $\psi(\rho=0)=\psi^{\prime}(\rho)|_{\rho=0}=0$ Xiong et al. (2023). The vortex density reads
$n_{+}(\rho)=\exp(-\beta\psi+\gamma_{+})=\frac{4e^{\gamma_{+}}}{(\beta\rho^{2}e^{\gamma_{+}}-2)^{2}}.$ (72)
@@
[…]
which gives rise the relation between $e^{\gamma_{+}}$ and $\beta$ [Fig. 10(a) in Appendix]. When $\beta\rightarrow\beta_{s}=-16\pi$, vortices around the pole collapse into a point, known as the supercondensation. Hence the parameter range for the sandwich state is $\beta_{s}<\beta<\beta_{c}$. The limit configuration at $\beta=\beta_{s}$ is $n_{+}(\theta)=[\delta(\theta)+\delta(\theta-\pi)]/(4\pi\sin\theta)$. It is worthwhile mentioning that the value of $\beta_{s}$ depends on the normalization condition. For vortices confined to a disc on a plane, the vortex density profile is the same as Eq. (72), however the vortex number is normalized to unity, which gives rise that $\beta_{s}=-8\pi$ (in convention of this paper) Smith and O’Neil (1990); Yu et al. (2016); Kiessling (1995); Patil and Dunkel (2021); Xiong et al. (2023).
In spherical coordinates, the stream function and the vortices density on the southern hemispheres [$\theta\in(\pi/2,\pi)$] are [see also Fig. 3(a)]
$\psi(\theta)$ $=-\frac{2}{\beta}\log\left[\frac{2}{\beta\cot^{2}\left(\frac{\theta}{2}\right)e^{\gamma_{+}}-2}\right],$ (74)
$n_{+}(\theta)$ $=\frac{4e^{\gamma_{+}}}{\left[\beta\cot^{2}\left(\frac{\theta}{2}\right)e^{\gamma_{+}}-2\right]^{2}}.$ (75)
@@
Now we consider the distribution of negative vortices in the region near the equator which is mapped to the region near the edge of the disc. For $|z|\sim 1$, Eq. (68) can be well approximated as
$4\partial_{z}\partial_{\bar{z}}\psi$ $=e^{\beta\psi(z,\bar{z})+\gamma_{-}}.$ (76)
@@
$\frac{1}{\rho}\frac{{\rm d}}{{\rm d}\rho}\left(\rho\frac{{\rm d}}{{\rm d}\rho}\psi\right)$ $=e^{\beta\psi(\rho)+\gamma_{-}},$ (77)
and admits an exact solution
@@
$\psi(\rho=1)=\psi^{\prime}(\rho)|_{\rho=1}=0.$ (79)
Here $A=\pm\sqrt{4-2\beta e^{\gamma_{-}}}$. As it follows, the density is
$n_{-}(\rho)=\frac{4A^{2}\rho^{(A-2)}e^{\gamma_{-}}}{\left[2+A+(A-2)\rho^{A}\right]^{2}}.$ (80)
The relation between $e^{\gamma_{-}}$ and $\beta$ is given by the normalization condition Eq. (73) ($n_{+}\rightarrow n_{-}$). Moreover, the smooth condition $n^{\prime}_{-}(\rho)|_{\rho=0}=0$ requires $e^{\gamma_{-}}\beta<-5/2$ which gives rise to that $\beta<-7.11\pi$ [Fig. 10 (b) in Appendix]. Hence for $\beta$ within the range $\beta_{s}<\beta<\beta_{c}$, the smooth condition is automatically satisfied.
In spherical coordinates, the stream function and the vortex density distribution of negative vortices near the equator [$\theta\in(\pi/2,\pi)$] read
@@
$n_{-}(\theta)$ $=\frac{4A^{2}\cot^{(A-2)}(\theta/2)e^{\gamma_{-}}}{\left[2+A+(A-2)\cot^{A}(\theta/2)\right]^{2}}.$ (82)
Induced by the boundary condition Eq. (79), the fluid velocity $u_{\phi}=-\partial\psi/\partial\theta$ on the equator vanishes, namely $u_{\phi}(\theta=\pi/2)=\psi^{\prime}(\rho)\rho^{\prime}(\theta)|_{\theta=\pi/2}=0$, which is consistent with the fluid velocity field on the equator at low energies [$\psi_{20}^{\prime}(\theta)|_{\theta=\pi/2}=0$]. At $\beta=\beta_{s}$, $e^{\gamma_{-}}\approx 0.167,A\approx 4.556$ [Fig. 10(b) in Appendix], the vorticity for the sandwich state reads $\sigma_{s}(\textbf{r})=\Big{[}\delta(\theta)+\delta(\theta-\pi)\Big{]}/(4\pi\sin\theta)-n_{-}(\theta),$ and the quadrupole moment reaches the maximum value $Q^{\rm max}_{s}\approx 1.59$. It is easy to see that the quadrupole state also reaches the high energy limit at $\beta=\beta_{s}$, and the corresponding vorticity is ${\sigma}_{q}(\textbf{r})=\Big{[}\delta(\theta-\pi/4,\phi)+\delta(\theta-3\pi/4,\phi-\pi)-\delta(\theta-3\pi/4,\phi)-\delta(\theta-\pi/4,\phi-\pi)\Big{]}/(2\sin\theta)$ which gives rise the maximum quadrupole moment $Q^{\rm max}_{q}=3\sqrt{2}/2>Q^{\rm max}_{s}$. Note that for the quadrupole state locations of the clusters at supercondensation coincide with mechanical equilibrium positions for $4$ point vortices on a sphere.
IV.3.2 Dipole states at high energies
At high energies, for the dipole state, vortices with opposite sign are concentrated near the poles. The density distribution of positive vortices near the southern pole is identical to Eq. (72) and the only difference is that here the vortex density is normalized to unity but not 1/2. Hence for the dipole state, the supercondensation occurs at $\beta=\beta^{d}_{s}=-8\pi$ which is the same as what for vortices confined to a disc on a plane Smith and O’Neil (1990); Yu et al. (2016); Xiong et al. (2023). In spherical coordinates, vortex densities are [see also Fig. 4]
$n_{-}(\theta)$ $=\frac{4e^{\gamma_{-}}}{\Big{[}\beta\cot^{2}(\theta/2)e^{\gamma_{-}}-2\Big{]}^{2}}\quad\theta\in(\pi/2,\pi),$ (83)
$n_{+}(\theta)$ $=\frac{4e^{\gamma_{+}}}{\Big{[}\beta\tan^{2}(\theta/2)e^{\gamma_{+}}-2\Big{]}^{2}}\quad\theta\in(0,\pi/2),$ (84)
@@ IV.3.2 Dipole states at high energies
$L_{\zeta}(\beta)=$ $\frac{8\pi e^{\gamma_{-}}}{(2+\beta e^{\gamma_{-}})^{4}}\Big{[}\beta e^{\gamma_{-}}\Big{(}\beta e^{\gamma_{-}}(-2\beta+5+32\log 2)$ (85)
$+20-32\log 2\Big{)}-16\beta e^{\gamma_{-}}(\beta e^{\gamma_{-}}-1)\log(2-\beta e^{\gamma_{-}})+4\Big{]}.$
@@
When $\beta\rightarrow\beta^{d}_{s}$, $n_{-}=\delta(\theta-\pi)/(2\pi\sin\theta),n_{+}=\delta(\theta)/(2\pi\sin\theta)$, and the angular momentum reaches the upper bound $L_{\zeta}=L^{\rm{max}}_{\zeta}=2$.
For the dipole state at low energies, the fluid velocity $u_{\theta}=0$, $u_{\phi}\propto\sin\theta$ and $u_{\phi}$ vanishes at poles (Fig. 5). At high energies, the velocity field is
$u_{\phi}(\theta)=-\frac{\partial\psi}{\partial\theta}=-\frac{2\cot(\theta/2)\gamma_{-}}{[-2+\beta\gamma_{-}\cot^{2}(\theta/2)]\sin^{2}(\theta/2)},$ (86)
@@
Figure 5: The velocity field $u_{\phi}$ for the dipole state at a low energy (orange solid line) and at a high energy (blue solid line). The parameter region for the dipole state is $-8\pi=\beta^{d}_{s}<\beta<\beta^{d}_{c}=-4\pi$. When $\beta\to\beta^{d}_{s}$, $u_{\phi}(\theta=\pi/2)\rightarrow 1/2\pi$. Note that here the fluid velocity field on the northern hemisphere is obtained by shifting the coordinate of Eq. (86) ($\theta\rightarrow\pi-\theta$). The two expressions work well only near the poles and take different values on the equator. However the difference decreases as increasing energy and for the $\beta$ value we choose here, the discrepancy is invisible.
The quadrupole state at high energies can be easily constructed from Eqs. (83) and  (84). From the approximate high energy expressions of vortex densities, we can calculate the entropy and find again that $S[\text{sandwich state}]>S[\text{quadrupole state}]$.
IV.3.3 General solutions
A general solution to Eq. (69) is available Crowdy (1997); Henrici (1993) :
$\psi(z,\bar{z})=$ $-\frac{1}{\beta}\log\left[\frac{2f^{\prime}(z)\bar{f}^{\prime}(\bar{z})}{-c\beta(1+f(z)\bar{f}(\bar{z}))^{2}}\right],$ (87)
$n(z,\bar{z})=$ $c\exp(-\beta\psi),$ (88)
where $f(z)$ is a meromorphic function of $z$ in a simply connected domain and $\bar{f}(\bar{z})$ is the complex conjugate of $f(z)$. Furthermore , it is required that $f(z)$ has at most simple poles and $f^{\prime}(z)\neq 0$ in the consider domain Henrici (1993); CROWDY (2004).
Here we show that Eqs. (71) and (78) fall into this general solution by specifying function $f(z)$ properly. Let us firstly consider the vortex distribution around the pole. For an axisymmetric distribution, the stream function $\psi(z,\bar{z})$ must be the function of $z\bar{z}$ or $\rho$ only. Hence $f(z)$ has a unique form and $f(z)=az^{b}$, here $a\in\mathbb{C},b\in\mathbb{Z}$. The boundary condition $\psi(\rho=0)=\psi^{\prime}(\rho)|_{\rho=0}=0$ determines that $a=\sqrt{-c\beta/2}$, $b=1$, and $c=e^{\gamma_{+}}$, which leads to  Eq. (71) precisely.
For the distribution near the edge (or equator in spherical coordinates), the general solution Eq. (87) gives rise to Eq. (78) by choosing $f(z)=\ \sqrt{(A-2)/(A+2)}\ z^{A/2}$, where $A=\pm\sqrt{4-2\beta e^{\gamma_{-}}}$, and $c=e^{\gamma_{-}}/4$. Here the boundary condition $\psi(\rho=1)=\psi^{\prime}(\rho)|_{\rho=1}=0$ is fulfilled. However it is important to notice that the exponent $A/2$ is not an integer in general. Such a choice of $f(z)$ may not be a meromorphic function and it does not meet the requirements for Eq. (87). As shown previously, such a solution is indeed the solution to Eq. (77), suggesting that the validity condition of Eq. (88) can be extended.
IV.4 Gaussian clustered states
@@
$\psi(\theta)=$ $-\frac{1}{8\pi\sinh\alpha}\Big{[}e^{\alpha}{\rm Ei}\Big{(}\alpha(\cos\theta-1)\Big{)}-e^{\alpha}{\rm Ei}\Big{(}-\alpha(\cos\theta+1)\Big{)}$
$+e^{-\alpha}{\rm Ei}\Big{(}\alpha(1-\cos\theta)\Big{)}-e^{-\alpha}{\rm Ei}\Big{(}\alpha(\cos\theta+1)\Big{)}$
@@
where $\alpha\in(-\infty,\infty)$, ${\rm Ei}(x)=\int^{x}_{-\infty}e^{t}/t\ {\rm d}t$, and $C$ is a constant determined by proper smooth conditions. The angular momentum for this Gaussian state is
@@
The fluid velocity field for this state is
$u_{\theta}$ $=\frac{1}{\sin\theta}\frac{\partial\psi}{\partial\phi}=0,$ (92)
@@
Here we impose the zero velocity condition at poles, namely $u_{\phi}(\theta=0)=u_{\phi}(\theta=\pi)=0$, giving rise that $C=-2\cosh(\alpha)/\alpha$. Figure 6 shows the vorticity distribution, the energy and the angular momentum as functions of $\alpha$. Note that the condition $u_{\phi}(\theta=\pi/2)=0$ leads to unphysical results $u_{\phi}(\theta=0)=u_{\phi}(\theta=\pi)=\infty$ for $\alpha\neq 0$.
@@
V Microcanonical Monte Carlo Simulations

### v2

@@ III.4.1 Sandwich states at high energies
$\frac{1}{4}\frac{1}{\rho}\frac{{\rm d}}{{\rm d}\rho}\left(\rho\frac{{\rm d}}{{\rm d}\rho}\psi\right)=-{\rm e}^{-\beta\psi(\rho)+\gamma_{+}},$ (69)
@@
$\psi(\rho)$ $=-\frac{2}{\beta}\log\left(\frac{2}{\beta\rho^{2}{\rm e}^{\gamma_{+}}-2}\right)$ (70)
with boundary condition $\psi(\rho=0)=\psi^{\prime}(\rho)|_{\rho=0}=0$ Smith and O’Neil (1990); Kiessling (1995); Patil and Dunkel (2021); Xiong et al. (2023). Here the boundary condition $\psi^{\prime}(\rho)|_{\rho=0}=0$ for the problem of chiral vortices confined to a disc is induced by the flow generated by the vortices on the sphere and can be obtained from the global stream function Eq. (54) at low energies. Also the vortex density distribution Eq. (71) does not depend on the value of $\psi(\rho=0)$ and here we choose $\psi(\rho=0)=0$ for convenience [see Sec. B in Appendix]. We emphasize again that the exact solution Eq. (70) is valid in the whole disc while provides a good approximation of the distribution of vortices near the poles at large energies.
The vortex density then reads
$n_{+}(\rho)=\exp(-\beta\psi+\gamma_{+})=\frac{4{\rm e}^{\gamma_{+}}}{(\beta\rho^{2}{\rm e}^{\gamma_{+}}-2)^{2}}.$ (71)
@@
$\frac{1}{2}=\int^{2\pi}_{0}{\rm d}\phi\int_{0}^{1}\frac{4\rho}{(1+\rho^{2})^{2}}\,{\rm d}\rho\,n_{+}(\rho),$ (72)
which gives rise to the relation between ${\rm e}^{\gamma_{+}}$ and $\beta$ [Fig. C(a) in Appendix].
When $\beta\rightarrow\beta_{s}=-16\pi$, vortices around the pole collapse into a point, known as the supercondensation. Hence the parameter range for the sandwich state is $\beta_{s}<\beta<\beta_{c}$. The limit configuration at $\beta=\beta_{s}$ is $n_{+}(\theta)=[\delta(\theta)+\delta(\theta-\pi)]/(4\pi\sin\theta)$. It is worthwhile mentioning that the value of $\beta_{s}$ depends on the normalization condition. For vortices confined to a disc on a plane, the vortex density profile is the same as Eq. (71), however the vortex number is normalized to unity, which gives rise to that $\beta_{s}=-8\pi$ (in convention of this paper) Smith and O’Neil (1990); Yu et al. (2016); Kiessling (1995); Patil and Dunkel (2021); Xiong et al. (2023).
In spherical coordinates, the stream function and the vortices density on the southern hemispheres $\theta\in(\pi/2,\pi)$] are [see also Fig. 3(a)]
$\psi(\theta)$ $=-\frac{2}{\beta}\log\left[\frac{2}{\beta\cot^{2}\left(\frac{\theta}{2}\right){\rm e}^{\gamma_{+}}-2}\right],$ (73)
$n_{+}(\theta)$ $=\frac{4{\rm e}^{\gamma_{+}}}{\left[\beta\cot^{2}\left(\frac{\theta}{2}\right){\rm e}^{\gamma_{+}}-2\right]^{2}}.$ (74)
@@
Now we consider the distribution of negative vortices in the region near the equator which is mapped to the region near the edge of the disc. For $|z|\sim 1$, Eq. (67) is well approximated as
$4\partial_{z}\partial_{\bar{z}}\psi$ $={\rm e}^{\beta\psi(z,\bar{z})+\gamma_{-}}.$ (75)
@@
$\frac{1}{\rho}\frac{{\rm d}}{{\rm d}\rho}\left(\rho\frac{{\rm d}}{{\rm d}\rho}\psi\right)$ $={\rm e}^{\beta\psi(\rho)+\gamma_{-}},$ (76)
which admits an exact solution
@@
$\psi(\rho=1)=\psi^{\prime}(\rho)|_{\rho=1}=0$ (78)
and $A=\pm\sqrt{4-2\beta{\rm e}^{\gamma_{-}}}$. Similarly, the exact solution Eq. (77) to the problem of chiral vortex clusters in a disc with the boundary condition Eq. (78) is well-defined on the whole disc domain and provides a good approximation of the distribution of vortices near the equator at large energies.
As it follows, the density is
$n_{-}(\rho)=\frac{4A^{2}\rho^{(A-2)}{\rm e}^{\gamma_{-}}}{\left[2+A+(A-2)\rho^{A}\right]^{2}}.$ (79)
Note that the boundary condition $\psi^{\prime}(\rho)|_{\rho=1}=0$ is also induced by the global stream function Eq. (54). The vortex density distribution Eq. (79) is independent of the value of $\psi(\rho=1)$ and here we choose $\psi(\rho=1)=0$ for convenience [see Sec. B in Appendix]. The relation between ${\rm e}^{\gamma_{-}}$ and $\beta$ is given by the normalization condition Eq. (72) ($n_{+}\rightarrow n_{-}$). Moreover, the smooth condition $n^{\prime}_{-}(\rho)|_{\rho=0}=0$ requires ${\rm e}^{\gamma_{-}}\beta<-5/2$ which gives rise to that $\beta<-7.11\pi$ [Fig. C (b) in Appendix]. Hence for $\beta$ within the range $\beta_{s}<\beta<\beta_{c}$, the smooth condition is automatically satisfied.
In spherical coordinates, the stream function and the vortex density distribution of negative vortices near the equator $\theta\in(\pi/2,\pi)$] read
@@
$n_{-}(\theta)$ $=\frac{4A^{2}\cot^{(A-2)}(\theta/2){\rm e}^{\gamma_{-}}}{\left[2+A+(A-2)\cot^{A}(\theta/2)\right]^{2}}.$ (81)
Induced by the boundary condition Eq. (78), the fluid velocity $u_{\phi}=-\partial\psi(\theta)/\partial\theta$ on the equator vanishes, namely $u_{\phi}(\theta=\pi/2)=0$, which is consistent with the fluid velocity field on the equator at low energies $\psi_{20}^{\prime}(\theta)|_{\theta=\pi/2}=0$]. At $\beta=\beta_{s}$, ${\rm e}^{\gamma_{-}}\approx 0.167,A\approx 4.556$ [Fig. C(b) in Appendix], the vorticity for the sandwich state reads $\sigma_{s}(\bm{r})=\Big[\delta(\theta)+\delta(\theta-\pi)\Big]\Big/(4\pi\sin\theta)-n_{-}(\theta),$ and the quadrupole moment reaches the maximum value $Q^{\rm max}_{s}\approx 1.59$. It is easy to see that the quadrupole state also reaches the high energy limit at $\beta=\beta_{s}$, and the corresponding vorticity is ${\sigma}_{q}(\bm{r})=\Big[\delta(\theta-\pi/4,\phi)+\delta(\theta-3\pi/4,\phi-\pi)-\delta(\theta-3\pi/4,\phi)-\delta(\theta-\pi/4,\phi-\pi)\Big]/(2\sin\theta)$ which gives rise to the maximum quadrupole moment $Q^{\rm max}_{q}=3\sqrt{2}/2>Q^{\rm max}_{s}$. Note that for the quadrupole state locations of the clusters at supercondensation coincide with mechanical equilibrium positions for $4$ point vortices on a sphere.
III.4.2 General solutions
A general solution to Eq. (68) is available Crowdy (1997); Henrici (1993) :
$\psi(z,\bar{z})=$ $-\frac{1}{\beta}\log\left[\frac{2f^{\prime}(z)\overline{f^{\prime}(z)}}{-c\beta[1+f(z)\overline{f(z)}]^{2}}\right],$ (82)
$n(z,\bar{z})=$ $c\exp(-\beta\psi),$ (83)
where $f(z)$ is a meromorphic function of $z$ in a simply connected domain and $\overline{f(z)}$ is the complex conjugate of $f(z)$. Furthermore, it is required that $f(z)$ has at most simple poles and $f^{\prime}(z)\neq 0$ in the consider domain Henrici (1993); CROWDY (2004).
[…]
III.4.3 Dipole states at high energies
At high energies, for the dipole state, vortices with opposite sign are concentrated near the poles. The density distribution of positive vortices near the southern pole is identical to Eq. (71) and the only difference is that here the vortex density is normalized to unity but not 1/2. Hence for the dipole state, the supercondensation occurs at $\beta=\beta^{d}_{s}=-8\pi$ which is the same as what for vortices confined to a disc on a plane Smith and O’Neil (1990); Yu et al. (2016); Xiong et al. (2023).
In spherical coordinates, vortex densities are [see also Fig. 4]
$n_{-}(\theta)$ $=\frac{4{\rm e}^{\gamma_{-}}}{\Big[\beta\cot^{2}(\theta/2){\rm e}^{\gamma_{-}}-2\Big]^{2}}\quad\theta\in(\pi/2,\pi),$ (84)
$n_{+}(\theta)$ $=\frac{4{\rm e}^{\gamma_{+}}}{\Big[\beta\tan^{2}(\theta/2){\rm e}^{\gamma_{+}}-2\Big]^{2}}\quad\theta\in(0,\pi/2),$ (85)
@@ III.4.3 Dipole states at high energies
$L_{\zeta}(\beta)=$ $\frac{8\pi{\rm e}^{\gamma_{-}}}{(2+\beta{\rm e}^{\gamma_{-}})^{4}}\Big[\beta{\rm e}^{\gamma_{-}}\Big(\beta{\rm e}^{\gamma_{-}}(-2\beta+5+32\log 2)$ (86)
$+20-32\log 2\Big)-16\beta{\rm e}^{\gamma_{-}}(\beta{\rm e}^{\gamma_{-}}-1)\log(2-\beta{\rm e}^{\gamma_{-}})+4\Big].$
When $\beta\rightarrow\beta^{d}_{s}$, $n_{-}\rightarrow\delta(\theta-\pi)/(2\pi\sin\theta),n_{+}\rightarrow\delta(\theta)/(2\pi\sin\theta)$, and the angular momentum reaches the upper bound $L_{\zeta}\rightarrow L^{\rm{max}}_{\zeta}=2$.
@@
We now consider the flow generated by these vortices. The fluid velocity field reads $\bm{u}=\nabla\psi\times\bm{e}_{r}=u_{\theta}\bm{e}_{\theta}+u_{\phi}\bm{e}_{\phi}$, where $\bm{e}_{r}$, $\bm{e}_{\theta}$, $\bm{e}_{\phi}$ are three orthogonal spherical unit vectors. For the dipole state at low energies, $u_{\theta}=(1/\sin\theta)\partial\psi/\partial\phi=0$, $u_{\phi}=-\partial\psi/\partial\theta\propto\sin\theta$ and $u_{\phi}$ vanishes at poles (Fig. 5). Here At high energies, the velocity field is
$u_{\phi}(\theta)=-\frac{\partial\psi}{\partial\theta}=\frac{2\cot(\theta/2)\gamma_{-}}{[2-\beta\gamma_{-}\cot^{2}(\theta/2)]\sin^{2}(\theta/2)},$ (87)
@@
Figure 5: The velocity field $u_{\phi}$ for the dipole state at a low energy (orange solid line) and at a high energy (blue solid line). The parameter region for the dipole state is $8\pi=\beta^{d}_{s}<\beta<\beta^{d}_{c}=-4\pi$. When $\beta\to\beta^{d}_{s}$, $u_{\phi}(\theta=\pi/2)\rightarrow 1/2\pi$. Note that here the fluid velocity field on the northern hemisphere is obtained by shifting the coordinate of Eq. (87) ($\theta\rightarrow\pi-\theta$). The two expressions work well only near the poles and take different values on the equator. However the difference decreases as increasing energy and for the $\beta$ value we choose here, the discrepancy is invisible.
III.5 Comparison of the entropy between the sandwich state and the quadrupole state at high energies
The quadrupole state at high energies can be easily constructed from Eqs. (84) and  (85). From the approximate high energy expressions of vortex densities, we can calculate the entropy and find again that S[sandwich state]>S[quadrupole state]S[$\text${sandwich state}]>S[$\text${quadrupole state}] (Fig. 6).
Figure 6: A comparison between the entropy of the sandwich state (red circle) and the quadrupole state (blue cross) at high energies. The entropy values are evaluated using Eqs. (74) and  (81) for the sandwich state and Eqs. (84) and (85) for the quadrupole state.
III.6 Gaussian clustered states
@@
$\psi(\theta)=$ $-\frac{1}{8\pi\sinh\alpha}\Big[{\rm e}^{\alpha}{\rm Ei}\Big(\alpha(\cos\theta-1)\Big)-{\rm e}^{\alpha}{\rm Ei}\Big(-\alpha(\cos\theta+1)\Big)$
$+{\rm e}^{-\alpha}{\rm Ei}\Big(\alpha(1-\cos\theta)\Big)-{\rm e}^{-\alpha}{\rm Ei}\Big(\alpha(\cos\theta+1)\Big)$
@@
where $\alpha\in(-\infty,\infty)$, ${\rm Ei}(x)=\int^{x}_{-\infty}{\rm e}^{t}/t\ {\rm d}t$, and $C$ is a constant determined by proper smooth conditions. The angular momentum for this Gaussian state is
@@
The components of the fluid velocity field for this state are $u_{\theta}=0$ and
@@
Here we impose the zero velocity condition at poles, namely $u_{\phi}(\theta=0)=u_{\phi}(\theta=\pi)=0$, giving rise to that $C=-2\cosh(\alpha)/\alpha$. Figure 7 shows the vorticity distribution, the energy and the angular momentum as functions of $\alpha$. Note that the condition $u_{\phi}(\theta=\pi/2)=0$ leads to unphysical results $u_{\phi}(\theta=0)=u_{\phi}(\theta=\pi)=\infty$ for $\alpha\neq 0$.
@@
IV Monte Carlo Simulations